OpenStudy (anonymous):

The attendance last year at the Rock Brothers Music Festival was 567 people. This year the attendance was 670 people. What is the percent of increase for this year? Round to the nearest whole percent. A.18% B.20% C.15% D.19%

OpenStudy (anonymous):

what is the increase anna?

OpenStudy (anonymous):

ont typ the answer harkirat

OpenStudy (anonymous):

it is 18.17 i.e 18%

OpenStudy (anonymous):

sorry i had pressed the button

OpenStudy (anonymous):

Thanks. (:

OpenStudy (anonymous):

Anna you can get the answer as follows : first calculate the increase i.e 670 -567 = 103 (increase/attendance for last year) x 100% gives the answer

OpenStudy (anonymous):

% increase= (670-567)*100/567=18%

OpenStudy (anonymous):

pls click GOOD ANS button for me :))